OpenStudy (anonymous):

Simplify the expression. 3^7/3^5

jimthompson5910 (jim_thompson5910):

when you divide exponential expressions with the same base, you subtract the exponents

jimthompson5910 (jim_thompson5910):

so 3^7/3^5 turns into 3^(7-5) which then becomes 3^2 Finally, 3^2 = 3*3 = 9 So 3^7/3^5 = 9